Bend Ski and Board Shop sell snowboards. They made the following purchases of the Supermodel X Free-Rider Board from Burton Co. during the year:

20 at $115 in Sept

40 at $120 in Oct;

20 at $125 in Nov

By the end of the year, 64 snowboards were sold.

The Weighted Average Cost of Goods Sold is:

a. $ 7,760

b.$ 7,680

c.$ 7,600
Answer #1

Ending Inventory = 20+40+20 - 64 = 16 units

Weighted Average Cost per unit = (20*115+40*120+20*125)/(20+40+20) = $120 per unit

Weighted Average Cost of Goods Sold = 64*120 = $7680

Option b. is correct answer.